Two wires are made of the same material and have the same volume. However wire 1 has cross-sectional area A and wire 2 has cross-sectional area 3A. if the length of wire 1 increases by ?x on applying force F, how much force is needed to stretch wire 2 by the same amount ?

Grade:11

2 Answers

Arun
3 years ago

Since the two wires have same volume.
Length of wire 1 =l=l
Length of wire 2 =l3=l3
Area of wire 1 =A=A
Area of wire 2 =3A=3A
wire 1:
Y=F/AΔx/lY=F/AΔx/l---------(1)
wire 2:
Y=F′/AΔx/l/3Y=F′/AΔx/l/3-------------(2)
from (1) and (2)
FA×lΔx=F13A×l3ΔxFA×lΔx=F13A×l3Δx
F1=9FF1=9F
 
Vikas TU
3 years ago
Length of wire 1 =l=l
Length of wire 2 =l3=l3
Area of wire 1 =A=A
Area of wire 2 =3A=3A
wire 1:
Y=F/AΔx/lY=F/AΔx/l---------(1)
wire 2:
Y=F′/AΔx/l/3Y=F′/AΔx/l/3-------------(2)
from (1) and (2)
FA×lΔx=F13A×l3ΔxFA×lΔx=F13A×l3Δx
F1=9FF1=9F